Management Meeting

2019-12-20 12:00
America/New York
Friday, 20 December 2019
, at 17:00 (GMT), duration : 00:00
BlueJeans Conference ID : 160455989

Agenda:
  • News from the run
    • Going reasonably well. Luminosity close to factor 2 lower than hoped. Being investigated, if stays like this will seriously affect collected data stats. CAD say beam at max of what they can do, but luminosity low.
    • Test of FXT going to happen over the weekend, eTOF going well
    • No other issues. Shifts seem covered for now
  • Production
    • ToF calibration for FXT almost there. Last step before production starts
    • Run 17 some progress. Slow but steady
    • FMS Calibration
  • Embedding
    • Mismatch for pp 200. Progress using different libraries. Investigation continues- see S&C reports
    • Probably need to pause submissions/approvals using SL16d_embed or later.
  • New BulkCorr Convener
  • CoC
  • HLT discussion
  • Meeting time - new semester?
  • AOB

Management Meeting

2019-12-27 12:00
America/New York
Friday, 27 December 2019
, at 17:00 (GMT), duration : 00:00

Management Meeting

2020-01-03 12:00
America/New York
Friday, 3 January 2020
, at 17:00 (GMT), duration : 00:00

Management Meeting

2019-12-06 12:00
America/New York
Friday, 6 December 2019
, at 17:00 (GMT), duration : 00:00


BlueJeans Conference ID : 160455989,

Agenda
  • Run-20
  • Production and embedding update/questions
  • Celebrating 20 years at BNL in July
  • Collaboration meeting LBNL
  • Isobar 
    • ABC paper submission
  • pp 500
  • New Institute Application
  • New STC members
  • New HF convener
  • AOB

Management Meeting

2019-12-13 12:00
America/New York
Friday, 13 December 2019
, at 17:00 (GMT), duration : 00:00

Agenda:
  • Run report
    • Collisions now underway. Intensity (as of Wed) still a little low. 14 min fills likely operation mode.
    • TPC Lasers were fixed
    • eTOF data now in QA plots. Still some issues with DAQ/T0 being investigated. Stability of operations ongoing, progress mass
    • Other detectors coming up well
    • Weekly QA meeting started. New tool almost ready to automate the QA plots for the PWG - a summer project of Ben and an summer high school student at BNL. HLT and Fast Offline ongoing.
    • QA fast of Shift need to start a week earlier than planned. This is happening and experienced folks identified to take on the task as it ramps up.
  •  Production
    • Run 18: FxT. ToF calibration ongoing.
    • Run 19: TPC timing inner vs outer
    • Run 17: Production calibration samples. Look better but studies still ongoing, not yet perfect. Need to understand how good it needs to be for analyses planned.
  • Embedding
    • Run 16-18 TPC simulation of clusters not accurate enough due to ADC threshold change. Work ongoing.
    • Run 15 pp200. Data and simulation don't match. Material from HFT seems to be correct. Rather maybe GEANT run incorrectly. Being investigated.
    • Large net-proton embedding request submitted for 27 GeV. This is the PAC request to check non-binomial correction is correct

Management Meeting

2019-11-29 12:00
America/New York
Friday, 29 November 2019
, at 17:00 (GMT), duration : 00:00

Management Meeting

2019-11-08 12:00
America/New York
Friday, 8 November 2019
, at 17:00 (GMT), duration : 00:00

Management Meeting

2019-11-15 12:00
America/New York
Friday, 15 November 2019
, at 17:00 (GMT), duration : 00:00
 BlueJeans Conference ID : 160455989,

Agenda
  • Preparations for Run-20
    • Pole tip closed, interlock testing underway.
    • Trigger board setup and Chair? - Elke will ask JH to call for first meeting next week
    • RHIC coordination meeting on the 19th someone needs to be there.
  • Production and embedding update/questions
    • FXT calibrations
    • Cu+Cu - eta asymmetry
    • pp 500 calibration issues?
    • PicoDst reproduction for EMC due to 700 MeV cut
    • EMC in Isobar data
    • Embedding NHits still not perfect
    • Embedding for Isobar
    • PWGC rewrite embedding and production priority list
    • Embedding for 200 GeV central data for net-proton analysis
    • Need a list of what runs have what issues
  • Celebrating 20 years at BNL or LBNL?
  • News from QM:
    • v3 in small systems, request from BNL to provide analysis note clearly documenting the differences in measured v2 and v3 in the 3 systems between STAR and PHENIX. The document should clearly lay out where are the differences, and how significant are they. Step 2: we will attempt to get STAR and PHENIX together to investigate the origin of the discrepancies. This would preferably include groups that are members of both collaborations. To the extent possible, the investigation should attempt to select common acceptance. If not possible, we should attempt to get constraints on possible acceptance dependence from models.
    • 200 GeV d,p,t review - anomalously small value?
    • 200 GeV Tch vs Mu_B result
    • Isobar - presentation on expectations from simulations of signals,
      • ABC paper submission
    • 62.4 GeV Lambda results?
    • Large stats. net-proton embedding ; submission of techniques paper
  • Request  to add one name to the RP papers from Run 15. Stephen Bueltmann, who is at ODU is not a STAR collaborator anymore. Despite that he came to BNL in preparation for Run 15 and helped us install and commission Roman Pots at their new location between DX-D0 magnets. He was a key person on PP2PP and his expertise was essential in our smooth installation and commissioning for run 15.
    • Agreed this should happen, only question about how to list his Institute
  • AOB

Management Meeting

2019-11-22 12:00
America/New York
Friday, 22 November 2019
, at 17:00 (GMT), duration : 00:00
 BlueJeans Conference ID : 160455989,

Agenda
  • Preparations for Run-20
  • Production and embedding update/questions
  • pp 500 calilbration meeting
  • Isobar analysis
  • UTA's application
  • AOB

Management Meeting

2019-10-11 12:00
America/New York
Friday, 11 October 2019
, at 16:00 (GMT), duration : 00:00

BlueJeans Conference ID : 160455989, 
  • Summer Shut down/start of run
  • DNP
  • (pre-)QM meeting organization
  • Production, calibration, embedding
  • Isobar analysis paper
  • University of Tarapaca (UTA) - SOnja Kabana - applicication
  • T-Shirts have arrived

Management Meeting

2019-10-18 12:00
America/New York
Friday, 18 October 2019
, at 16:00 (GMT), duration : 00:00
BlueJeans Conference ID : 160455989,
  • Summer Shut down/start of run/PC sign-up
    • First forward upgrade detector installed
    • eTOF all installed as of this morning
    • PC signup now on shift signup
  • pre-QM meeting organization
    • Reviewers of posters and talks
  • post-QM meeting
    • What's the one topic - chiral, thermal and vortical properties of QCD. Function of rapidity
  • Production, calibration, embedding
    • pp  Run 17 - no new insights this week. People still looking, can we get people together at BNL soon
    • FXT - dEdx and TOF calibration checks underway
    • Run 19 - space charge calibration almost done. Will then launch TOF and dEdx calibration
    • nHits in embedding. Two comparison embeddings done for Run 18 Au-Au. Dielectron check done reveals no big efficiency difference. But at large eta both have less NHits than data. Uses TOF matching in data though and not in embedding. p over all eta seems fine, working on differential comparison in eta
    • EMC-track matching issue in picoDsts. 700 MeV min Tower cut an issue.
  • Isobar analysis paper
    • Release to collab for comments next week. Submit to arXiv before QM.
  • AOB
    • Jan meeting on post BES-II and Cold QCD. Will ping key Cold QCD and forward upgrade people for dates.
    • We encourage considering the use of Mattermost the next step for communication. Good STAR is doing this.

Management Meeting

2019-10-25 12:00
America/New York
Friday, 25 October 2019
, at 16:00 (GMT), duration : 00:00
 

Management Meeting

2019-11-01 12:00
America/New York
Friday, 1 November 2019
, at 16:00 (GMT), duration : 00:00

Management Meeting

2019-09-27 12:00
America/New York
Friday, 27 September 2019
, at 16:00 (GMT), duration : 00:00

BlueJeans Conference ID : 160455989, at 16:00 (GMT), duration : 02:00

  • Summer shutdown status
    • EMCal installation waiting for HCal plate installation to finish. Potentially couple of days delay. Will be tight, needs extra hands to help with the installation
    • eTOF group arrive Oct 11
    • AC Unit shipping Oct 17th, should arrive end of Oct at BNL. Installation in Nov.
  • Run start and sign-up
    • Emails gone to Council members as to Authors - Due by end of September
    • Quota calculations underway
    • Period coordinators need to be identified
  • Productions and Embedding
    • Embedding going. Got 20% extra time on Cori
    • NHit still not resolved, proceeding with embedding with bug, keep working on fix
    • No fix/reason yet found for the pp h+/h- issue 
  • New DoE directive
    • Will ask specifically about does it apply to non-US visitors employed by non-US Institutions that  BNL group hosts
  • STAR Web manager
    • Looking for a replacement - volunteers welcome
  • LFSUPC and HF conveners
  • Update STC membership

Management Meeting

2019-10-04 12:00
America/New York
Friday, 4 October 2019
, at 16:00 (GMT), duration : 00:00


BlueJeans Conference ID : 160455989, 
  • Summer Shut down
  • DNP
  • pre-QM meeting organization
  • Shift organization
  • Production, calibration, embedding
  • Isobar analysis
  • AOB

Management Meeting

2019-09-13 12:00
America/New York
Friday, 13 September 2019
, at 16:00 (GMT), duration : 00:00
 
BlueJeans Conference ID : 160455989, at 16:00 (GMT), duration : 02:00
  • S&T review of STAR talks
  • Summer shutdown
  • Isobar status
  • Production and Embedding
  • QM and pre-QM meeting
  • Run Sign-up

Management Meeting

2019-09-20 12:00
America/New York
Friday, 20 September 2019
, at 16:00 (GMT), duration : 00:00

BlueJeans Conference ID : 160455989, at 16:00 (GMT), duration : 02:00
  1. Summer shutdown progress
  2. Embedding
  3. Productions
  4. S&T Review - Closeout summary
  5. AOB

Management Meeting

2019-08-30 12:00
America/New York
Friday, 30 August 2019
, at 16:00 (GMT), duration : 00:00

Management Meeting

2019-09-06 12:00
America/New York
Friday, 6 September 2019
, at 16:00 (GMT), duration : 00:00
  • Shutdown progress. So far going OK.
    • Shipping/handling issues at BNL - long delays causing potential issues for installation
    • Short shutdown means EEMC won't have time to do maintenance
  • EToF internal readiness review - Scheduled for Monday Sept. 16. Committee in place, agenda and charge almost complete
  • Production and Calibration:
    • Run-5 Cu+Cu started
    • Run-18 Fixed target. T0 of TPC calibration ongoing. Probably not ready for DNP, speaker(s) aware
    • Run-18 Au-Au 27: Still some issues with some bad runs, work ongoing to exclude them
    • Run-19 7.7: calibrations ongoing. 19.6: calibrations on space-charge, grid leak ongoing. New Temple student being trained.
    • Run-17 pp 510: Seems to be fill-to-fill issues. Hope to correct via space-charge/grid leak corrections
  • Embedding
    • Run-16 SL17dembed: Too few hits/track in the embedding compared to data. Fix still being worked on, if farm become free might try embedding with bug and then reproduce when fix made.
    • Run-18 Au-Au 27: iTPC sector row number simulation not matched to real data. Now fixed and embedding started

Management Meeting

2019-08-23 12:00
America/New York
Friday, 23 August 2019
, at 16:00 (GMT), duration : 00:00